# Frobenius algebra objects in Temperley-Lieb categories at roots of unity

### v2

@@ 2.1 Frobenius structures
An object $X\in\mathcal{C}$ is called invertible if it is left dualizable and $e_{X}$ and $c_{X}$ are both isomorphisms. Then $X^{\vee}$ is called the inverse object of $X$, and we have $X^{\vee}\cong{{}^{\vee}X}$.
@@
We have an isomorphism $e_{W}:V\otimes W\stackrel{{\scriptstyle\sim}}{{\to}}\mathds{1}$, so we construct an isomorphism as follows:
$f:V\otimes A\stackrel{{\scriptstyle 1_{V}\otimes\varphi}}{{\longrightarrow}}V\otimes W\otimes A^{\vee}\stackrel{{\scriptstyle e_{W}\otimes 1_{A^{\vee}}}}{{\longrightarrow}}A^{\vee}.$
@@
On composing with $e_{V}\otimes 1_{A}:W\otimes V\otimes A\to A$, applying the triangle identities shows that we get the identity map on $A$. So $e_{V}\otimes 1_{A}\circ\varphi_{\text{new}}=1_{A}$, and therefore $\varphi_{\text{new}}=(e_{V}\otimes 1_{A})^{-1}=c_{W}\otimes 1_{A}$. ∎
@@
    $A$ is left dualizable and $(A,u,m)$ has Frobenius structure $(W,n)$;
@@
[…]
@@ 2.2 Nakayama automorphisms
The category $\mathcal{C}$ is pivotal if there is an isomorphism of monoidal functors from the identity to $-^{\vee\vee}$: see [EGNO15, Section 4.7]. We say that an object $A\in\mathcal{C}$ is pivotal if the full subcategory on the object $A$ is pivotal.
@@ 2.3 Transfer of structure
Let $\mathcal{C}$, $\mathcal{D}$ be monoidal categories. Recall that a monoidal functor from $\mathcal{C}$ to $\mathcal{D}$ is a pair $(F,J)$ consisting of a functor $F:\mathcal{C}\to\mathcal{D}$ and isomorphisms $J_{C,C^{\prime}}:F(C)\otimes F(C^{\prime})\to(C\otimes C^{\prime})$ natural in both $C,C^{\prime}\in\mathcal{C}$. It should satisfy unitality and associativity conditions. Note that monoidal functors respect duals [EGNO15, Exercise 2.10.6].
@@
[…]
@@
Kepp the conditions and notation of Lemma 2.12. Applying the same methods as above, together with naturality of tensor products, gives the following result:
@@ 2.4 Examples
${\mathds{1}\otimes\mathds{1}\otimes\mathds{1}}$${\mathds{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathds{1}\otimes\mathds{1}\otimes X}$${X\otimes\mathds{1}\cong X}$${\mathds{1}\otimes X\otimes\mathds{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathds{1}\otimes X\otimes X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\otimes\mathds{1}\otimes\mathds{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\otimes\mathds{1}\otimes X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X\otimes X\cong\mathds{1}}$${X\otimes X\otimes\mathds{1}}$${X\otimes X\otimes X}$
@@
[…]
@@
$R$ is a $\beta$-Frobenius extension of $S$ if and only if there exists a map $n:R\to S_{\beta}$ such that $(S_{\alpha},n)$ is a Frobenius structure on $R$.
@@
First suppose that that $(S_{\alpha},n)$ is a Frobenius structure on $R$. As $R$ is left dualizable, we know that $R$ is finitely generated and projective as a right $S$-module. By assumption the map $\varphi:A\to W\otimes A^{\vee}$ from Definition 2.1 is an isomorphism $R\cong S_{\beta}\otimes_{B}\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S)$ of $S\text{-}S$-bimodules, and by Remark 2.2 it is in fact an isomorphism of $S\text{-}R$-bimodules. Finally, as $R$ is projective as a right $S$-module the natural map $S_{\beta}\otimes_{S}\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S)\to\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S_{\beta})$ of $S\text{-}R$-bimodules is an isomorphism.
@@ 3.1 Jones-Wenzl projections
  • Fix $k\geq 1$. Let $t=e^{\frac{\pi i}{2k+4}}$ be a fixed $42k+$ 4th root of unity and let $q=t^{2}=e^{\frac{\pi i}{k+2}}$.
  • Let $[i]=\frac{q^{i}-q^{-i}}{q-q^{-1}}\in\mathbb{R}$ denote the $i$th quantum integer and write $\delta=-[2]=-q-q^{-1}$.
@@
Note that $q^{k+2}=-1$, so $[k+2-j]=[j]$. In particular, $[k+2]=[0]=0$.
Let $1_{1}:1\to 1$, :02$\cup$:0$\to$ 2∪ : 0 → 2, and :20$\cap$:2$\to$ 0∩ : 2 → 0 denote the unique such morphisms in $\mathcal{D}$, then define
@@
both for $0\leq i\leq n$. Let $U_{i,n}=\cup_{i,n-2}\circ\cap_{i,n-2}:n\to n$ denote the cap-cup morphism for $21\leq i\leq n-$ 2, so $U_{i}^{2}=\delta U_{i}$ in $TL$. Define iterated cups $\Cup_{i,m,n}:n\to n+2m$ and caps $\Cap_{i,m,n}:n+2m\to n$ by
@@
Following [We87], after [Jo83], we have Jones-Wenzl projections $f_{i}:i\to i$ in $TL$, for $10\leq i\leq k+$ 1, defined recursively by $f_{0}=1_{0}$, $f_{1}=1_{1}$, and
@@
[…]
@@ 3.3 Twisted Frobenius structure
Then $\psi_{i}\circ\varphi_{i}=1_{F_{i}}$ by Proposition 3.1(v), using $[j]=[k+2-j]$ and $[1]=1$.
@@
$F_{i}$ $F_{k}$ $F_{k}$ $F_{i}$ $j$$i$$j$$=$ $F_{i}$ $F_{k}$ $F_{k}$ $F_{i}$ $j$$i$$j$
Now we use Lemma 3.18 twice to remove the cossings next to the $F_{k}$ boxes, which introduces a scalar $t^{2ij}$. Then we use Proposition 3.1 to remove the $F_{k}$ boxes, which introduces a scalar $1/[i+1]^{2}$, but this cancels with the scalar $[i+1]^{2}$ we omitted at the start of our calculation. So we have so far shown that $\alpha^{2}$ is the following map:
@@ 3.4 Preprojective algebras
The fact that the Nakayama automorphism of $\Pi$ has order $2$ can be used to show that the path algebra $\mathbb{C}Q$ of the Dynkin quiver $Q$ is fractionally Calabi-Yau: see [Gr23, Section 6.2.4]. This was originally proved by Miyachi and Yekutieli [MY01].

### v3

@@ 2.1 Frobenius structures
An object $X\in\mathcal{C}$ is called invertible if it is left dualizable and $e_{X}$ and $c_{X}$ are both isomorphisms. Then $X^{\vee}$ is called the inverse object of $X$. Note that the inverse of the triangle identities for $X$ show that $X^{\vee}$ has left dual $X$ with $e_{X^{\vee}}=c_{X}^{-1}$ and $c_{X^{\vee}}=e_{X}^{-1}$. Therefore ${{}^{\vee}X}\cong X^{\vee}$.
We now give some results on Frobenius structures $(W,n)$ under the assumption that $W$ is invertible.
@@
We have an isomorphism $e_{W}:V\otimes W\stackrel{{\scriptstyle\sim}}{{\to}}\mathds{1}$, so we construct a map
$f:V\otimes A\stackrel{{\scriptstyle 1_{V}\otimes\varphi}}{{\longrightarrow}}V\otimes W\otimes A^{\vee}\stackrel{{\scriptstyle e_{W}\otimes 1_{A^{\vee}}}}{{\longrightarrow}}A^{\vee}$
which, using $e_{W}^{-1}=c_{V}$, is an isomorphism with inverse
$f^{-1}:A^{\vee}\stackrel{{\scriptstyle c_{V}\otimes 1_{A^{\vee}}}}{{\longrightarrow}}V\otimes W\otimes A^{\vee}\stackrel{{\scriptstyle 1_{V}\otimes\varphi^{-1}}}{{\longrightarrow}}V\otimes A.$
@@
Using naturality of $\otimes$ and the triangle identities, we have:
$(e_{V}\otimes 1_{A})\circ\varphi_{\text{new}}$ $=(e_{V}\otimes 1_{A})\circ(1_{W}\otimes 1_{V}\otimes\varphi^{-1})\circ(1_{W}\otimes c_{V}\otimes 1_{A^{\vee}})\circ\varphi$
$=\varphi^{-1}\circ(e_{V}\otimes 1_{W}\otimes 1_{A^{\vee}})\circ(1_{W}\otimes c_{V}\otimes 1_{A^{\vee}})\circ\varphi$
$=\varphi^{-1}\circ(1_{W}\otimes 1_{A^{\vee}})\circ\varphi=1_{A}$
So $\varphi_{\text{new}}=(e_{V}\otimes 1_{A})^{-1}=c_{W}\otimes 1_{A}$. ∎
@@
    AA is left dualizable and $(A,u,m)$ has Frobenius structure $(W,n)$;
@@
The following diagrams illustrate the constructions used below:
[…]
@@
For $\emph{(ii)}\Rightarrow\emph{(i)}$ we can apply $\emph{(i)}\Rightarrow\emph{(ii)}$ in the monoidal category with opposite tensor product. ∎
@@ 2.2 Nakayama automorphisms
The category $\mathcal{C}$ is pivotal if there is an isomorphism of monoidal functors $\iota:$-$\to$-$^{\vee\vee}$ from the identity to the double dual: see [EGNO15, Section 4.7]. We say that an object $A\in\mathcal{C}$ is pivotal if the full subcategory on the object $A$ is pivotal.
@@ 2.3 Transfer of structure
Let $\mathcal{C}$, $\mathcal{D}$ be monoidal categories. Recall that a monoidal functor from $\mathcal{C}$ to $\mathcal{D}$ is a pair $(F,J)$ consisting of a functor $F:\mathcal{C}\to\mathcal{D}$ and isomorphisms $J_{C,C^{\prime}}:F(C)\otimes F(C^{\prime})\to F(C\otimes C^{\prime})$ natural in both $C,C^{\prime}\in\mathcal{C}$. It should satisfy unitality and associativity conditions. Note that monoidal functors respect duals [EGNO15, Exercise 2.10.6].
@@
${F(V\otimes A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F(\varphi_{A})}$${F(A\otimes V)}$${X\otimes B\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J_{V,A}}$$\scriptstyle{\varphi_{B}}$${B\otimes X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J_{A,V}}$
@@
Keep the conditions and notation of Lemma 2.12. Applying the same methods as above, together with naturality of tensor products, gives the following result:
@@ 2.4 Examples
[…]
@@
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a}$${2\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{b}$
@@
$R$ is a $\beta$-Frobenius extension of $S$ if and only if there exists a map $n:R\to S_{\beta}$ such that $(S_{\beta},n)$ is a Frobenius structure on $R$.
@@
First suppose that that $(S_{\beta},n)$ is a Frobenius structure on $R$. As $R$ is left dualizable, we know that $R$ is finitely generated and projective as a right $S$-module. By assumption the map $\varphi:A\to W\otimes A^{\vee}$ from Definition 2.1 is an isomorphism $R\cong S_{\beta}\otimes_{B}\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S)$ of $S\text{-}S$-bimodules, and by Remark 2.2 it is in fact an isomorphism of $S\text{-}R$-bimodules. Finally, as $R$ is projective as a right $S$-module the natural map $S_{\beta}\otimes_{S}\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S)\to\operatorname{Hom}\nolimits_{\operatorname{mod-}\!S}(R,S_{\beta})$ of $S\text{-}R$-bimodules is an isomorphism.
@@ 3.1 Jones-Wenzl projections
  • Fix $k\geq 1$. Let $t=e^{\frac{\pi i}{2k+4}}$ be a fixed $4k+8$th root of unity and let $q=t^{2}=e^{\frac{\pi i}{k+2}}$.
  • Let $i]=\frac{q^{i}-q^{-i}}{q-q^{-1}}\in\mathbb{R}$ denote the $i$th quantum integer and write $\delta=-[2]=-q-q^{-1}$.
@@
Note that $q^{k+2}=-1$, so [k+2j]=[j][k+2-j]=[j]. In particular, $[k+2]=[0]=0$.
Let $1_{1}:1\to 1$, $\cup:0\to 2$, and $\cap:2\to 0$ denote the unique such morphisms in $\mathcal{D}$, then define
@@
both for $0\leq i\leq n$. Let $U_{i,n}=\cup_{i,n-2}\circ\cap_{i,n-2}:n\to n$ denote the cap-cup morphism for $1\leq i\leq n-2$, so $U_{i}^{2}=\delta U_{i}$ in $TL$. Define iterated cups $\Cup_{i,m,n}:n\to n+2m$ and caps $\Cap_{i,m,n}:n+2m\to n$ by
@@
Following [We87], after [Jo83], we have Jones-Wenzl projections $f_{i}:i\to i$ in $TL$, for $0\leq i\leq k+1$, defined recursively by $f_{0}=1_{0}$, $f_{1}=1_{1}$, and
@@
${F_{h}\otimes F_{i}\otimes F_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes f_{i+j}}$$\scriptstyle{f_{h+i}\otimes 1}$${F_{h}\otimes F_{i+j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{h+i+j}}$${F_{h+i}\otimes F_{j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{h+i+j}}$${F_{h+i+j}}$
@@ 3.3 Twisted Frobenius structure
Then $\psi_{i}\circ\varphi_{i}=1_{F_{i}}$ by Proposition 3.1(v), using [j]=[k+2j][j]=[k+2-j] and $[1]=1$.
@@
$F_{i}$ $F_{k}$ $F_{k}$ $F_{i}$ jjiijj== $F_{i}$ $F_{k}$ $F_{k}$ $F_{i}$ jjiijj
Now we use Lemma 3.18 twice to remove the cossings next to the $F_{k}$ boxes, which introduces a scalar $t^{2ij}$. Then we use Proposition 3.1 to remove the $F_{k}$ boxes, which introduces a scalar $1/[i+1]^{2}$, but this cancels with the scalar $i+1]^{2}$ we omitted at the start of our calculation. So we have so far shown that $\alpha^{2}$ is the following map:
@@ 3.4 Preprojective algebras
The fact that the Nakayama automorphism of $\Pi$ has order 22 can be used to show that the path algebra $\mathbb{C}Q$ of the Dynkin quiver $Q$ is fractionally Calabi-Yau: see [Gr23, Section 6.2.4]. This was originally proved by Miyachi and Yekutieli [MY01].